OpenStudy (anonymous):

Donna ordered four burgers and one order of fries at the Hamburger Palace.However,the waiter put three burgers and two orders of fries in the bag and charged Donna the correct price for three burgers and two orders of fries,$3.15.When Donna discovered the mistake,she went back to complain.She found out the the price for four burgers and one order of fries is $3.15 and decided to keep what she had.What is the price of one burger,and what is the price of one order of fries?anyone can help :-(

OpenStudy (anonymous):

3b + 2f = 3.15 4b + f = 3.15 3b + 2f = 3.15 -2(4b + f = 3.15) 3b + 2f = 3.15 -8b - 2f = -6.30 -5b = -3.15 b = .63 f = .63 They both cost 63 cents

OpenStudy (anonymous):

wow thanks I really struggling this problem solving